OpenStudy (anonymous):

@ganeshie8 Which ordered pair is not a point on the graph of y = 1/2x - 7 ? A. (1, -6 ½) B. (-2, -8) C. (0, -7) D. (2, 8)

OpenStudy (solomonzelman):

Plug in your points into the equation.

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

how do i do -6 ½ ?

OpenStudy (anonymous):

-6.5 ?

OpenStudy (solomonzelman):

\(\large\color{blue}{y = 1/2x - 7 }\) for example last point, \(\large\color{blue}{8= 1/2~(2) - 7 }\) \(\large\color{blue}{8= 1 - 7 }\) \(\large\color{blue}{8= - 6 }\)

OpenStudy (solomonzelman):

So you see that the last option is not on the line.

OpenStudy (solomonzelman):

Do the rest of them like this.

OpenStudy (anonymous):

for the first one it would be -6.5 = ½(1) - 7 ?

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

ok ty

OpenStudy (solomonzelman):

yw

OpenStudy (solomonzelman):

but this is not the only point tho

OpenStudy (solomonzelman):

A B and C work

OpenStudy (solomonzelman):

only D doesn't

OpenStudy (anonymous):

and D doesn't yes i know :) thanks

OpenStudy (solomonzelman):

anytime